OpenStudy (anonymous):

I don't want the answer I would just like help doing it, Please! The graph of which equation has a vertex of (1, –3)? y = 1(x – 3)2 + 1 y = 2(x – 3)2 + 1 y = 2(x – 1)2 – 3 y = 2(x + 1)2 – 3

jimthompson5910 (jim_thompson5910):

Hint: y = a(x-h)^2 + k has the vertex of (h, k)

jimthompson5910 (jim_thompson5910):

So y = 1(x – 3)^2 + 1 for instance is in vertex form y = a(x-h)^2 + k where a = 1 h = 3 k = 1 So the vertex is (3, 1), which rules out choice A

OpenStudy (anonymous):

@jim_thompson5910 would the answer be D?

jimthompson5910 (jim_thompson5910):

y = 2(x + 1)^2 – 3 y = 2(x - (-1))^2 – 3 it's now in the form y = a(x-h)^2 + k where a = 2 h = -1 k = -3 So the vertex of this is (-1, -3)

jimthompson5910 (jim_thompson5910):

so close, just had the wrong sign for the x coordinate of the vertex

OpenStudy (anonymous):

so then it would be c?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

yw